Two point charges are 5 m apart and their
combined charge is 100 uC. If one repels
the other with a force of 0.9 N. Calculate
their values.
1
Expert's answer
2019-05-24T10:30:41-0400
"q_1 + q_2 = 100\cdot 10^{-6} C = 10^{-4} C"

On the other hand

"F = \frac{kq_1 q_2}{r^2}"

with k = 9 109 N m2/C2. The charges repel, so their values have the same sign. Hence


"\bigg\{\begin{matrix}
 q_1 + q_2 & = & 10^{-4} C \\
 q_1 q_2 & = & Fr^2/k = 25 \cdot 10^{-10} C^2 
\end{matrix}"




"(10^{-4} C - q_1) q_1 = 2.5 \cdot 10^{-9} C""q_1^2 -10^{-4} C \cdot q_1 + 2.5\cdot 10^{-9} C^2 = 0""q_1 = \frac{1}{2} \bigg( 10^{-4} \pm \sqrt{10^{-8} - 4\cdot 2.5 \cdot 10^{-9}}\bigg) C = 5 \cdot 10^{-5} C""q_2 = 5\cdot 10^{-5} C"